OpenStudy (anonymous):

evaluate the expression ,then convert your answer to decimal form . (2/3)^2=_____help plz

OpenStudy (hoblos):

(2/3)^2 =(2^2)/(3^2) = 4/9

OpenStudy (apoorvk):

 post ur answer and verify.